OpenStudy (anonymous):

5 sqrt 11 - 12 sqrt 11 - 2 sqrt 11

OpenStudy (anonymous):

since each term has sqrt 11 you can perform normal arithmetic

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

\[\LARGE 5\sqrt{11} - 12\sqrt{11} - 2\sqrt{11}\]

OpenStudy (anonymous):

yeah so you get -9sqrt(11)

OpenStudy (anonymous):

How'd you do that?

OpenStudy (anonymous):

(5-12-2)sqrt(11)